# June, 2024

### v1

@@ 2.2 Quantisation via Batalin-Vilkovisky formalism
$Z$ $=\int\big{[}\mathscr{D}\mbox{\boldmath$\varphi$}\big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\mbox{\boldmath$\varphi$}]\Big{)}~{},\qquad\mbox{\boldmath$\varphi$}=\Big{\{}\chi^{\alpha},\bar{\chi}_{\dot{\alpha}},V,C,\tilde{C},\eta,\bar{\eta},\mbox{\boldmath$\eta$},\bar{\mbox{\boldmath$\eta$}},\eta^{\prime},\bar{\eta}^{\prime}\Big{\}}~{},$
$S[\mbox{\boldmath$\varphi$}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}\frac{1}{8}\tilde{C}\Big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}+\frac{n}{2}\big{\{}\nabla^{2},\bar{\nabla}^{2}\big{\}}\Big{)}C+\bar{\eta}\mbox{\boldmath$\eta$}+\bar{\mbox{\boldmath$\eta$}}\eta$
$\qquad\qquad\qquad\qquad+m\bar{\eta}^{\prime}\eta^{\prime}-\frac{m}{8}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}^{2}\Big{\}}$
$\phantom{=}~{}-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,\big{[}\bar{\nabla}^{2}\nabla^{\alpha}\tilde{C},\bar{\nabla}^{2}\nabla_{\alpha}\tilde{C}\big{]}\eta+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\big{[}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\tilde{C},\nabla^{2}\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.28)
@@ 2.3 Quantum equivalence
$Z$ $=\int\big{[}\mathscr{D}V\big{]}\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}C\big{]}\big{[}\mathscr{D}\bar{\eta}\big{]}\big{[}\mathscr{D}\eta\big{]}\big{[}\mathscr{D}\bar{\mbox{\boldmath$\eta$}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\eta$}\big{]}\big{[}\mathscr{D}\pi\big{]}\delta(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi)\delta(\bar{W}_{\dot{\alpha}}-\frac{{\rm i}}{4}\nabla^{2}\bar{\nabla}_{\dot{\alpha}}\pi)$
×$\qquad\qquad\times\delta(\nabla^{2}\pi)\delta(\bar{\nabla}^{2}\pi)\text{Det}(\mathcal{H})\exp\Big{(}\frac{{\rm i}}{\hbar}\text{tr}\int{\rm d}^{4|4}z\,\Big{\{}V^{2}-2\tilde{C}\Box_{V}C+{\mbox{\boldmath$\eta$}}\bar{\eta}+{\eta}\bar{\mbox{\boldmath$\eta$}}\Big{\}}\Big{)}~{},$ (2.29)
@@
$\mathcal{H}$ $=\begin{pmatrix}0&-\frac{1}{4}\bar{\nabla}^{2}\\ -\frac{1}{4}\nabla^{2}&0\end{pmatrix}~{},$ (2.30a)
$\Box_{V}$ $=-\frac{1}{16}\Big{(}\nabla^{\alpha}\bar{\nabla}^{2}\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}\nabla^{2}\bar{\nabla}^{\dot{\alpha}}-\big{\{}\nabla^{2},\bar{\nabla}^{2}\big{\}}\Big{)}$
$=\nabla^{a}\nabla_{a}-\frac{1}{4}W^{\alpha}\nabla_{\alpha}+\frac{1}{4}\bar{W}_{\dot{\alpha}}\bar{\nabla}^{\dot{\alpha}}~{}.$ (2.30b)
The operator $\mathcal{H}$ acts on the space of covariantly chiral-antichiral pairs in the adjoint representation of $G$, while $\Box_{V}$ is defined to act on the space of unconstrained scalar superfields in the adjoint representation of $G$.
@@
$W_{\alpha}={\rm e}^{\bar{\Omega}_{Q}}\left(-\frac{1}{4}\bar{\cal D}^{2}\Big{(}{\rm e}^{-V_{Q}}{\cal D}_{\alpha}{\rm e}^{V_{Q}}\Big{)}+w_{\alpha}\right){\rm e}^{-\bar{\Omega}_{Q}}~{},$ (2.33)
@@
$W_{\alpha}\approx-\frac{1}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}V_{Q}+w_{\alpha}+\big{[}\bar{\Omega}_{Q},w_{\alpha}\big{]}~{}.$ (2.34)
@@
$w_{\alpha}=0\quad\implies\quad{\rm e}^{\bar{\omega}}={\rm e}^{\Phi}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.35)
In evaluating $\delta(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi)$, it suffices to keep only the terms of first order in the quantum prepotentials and $\pi$,
$\delta\left(W_{\alpha}+\frac{{\rm i}}{4}\bar{\nabla}^{2}\nabla_{\alpha}\pi\right)=\delta\left(-\frac{1}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}V_{Q}+\frac{{\rm i}}{4}\bar{\cal D}^{2}{\cal D}_{\alpha}\pi\right)~{}.$ (2.36)
@@
$V_{Q}$ $=\rho+\bar{\rho}+\frac{1}{2}\big{(}{\cal D}^{\alpha}\lambda_{\alpha}+\bar{{\cal D}}_{\dot{\alpha}}\bar{\lambda}^{\dot{\alpha}}\big{)}~{},\qquad\bar{{\cal D}}_{\dot{\alpha}}\rho=0~{},$ (2.37a)
@@
$1=\int\big{[}\mathscr{D}V_{Q}\big{]}\big{[}\mathscr{D}\pi\big{]}\exp\Big{(}\frac{{\rm i}}{2}\text{tr}\int\,{\rm d}^{4|4}z\,\Big{\{}{V_{Q}}^{2}-\pi^{2}\Big{\}}\Big{)}$ (2.38)
@@
~$\widetilde{\mathscr{H}}_{c}=\frac{1}{16}\bar{{\cal D}}^{2}{\cal D}^{2}~{},\qquad\widetilde{\mathscr{H}}_{a}=\frac{1}{16}{\cal D}^{2}\bar{{\cal D}}^{2}~{}.$ (2.40)
Here and below, a tilde over operators, such as $\widetilde{{\cal H}}$ and ~$\widetilde{{\square}}_{V}={\cal D}^{a}{\cal D}_{a}$, indicates that these operators are constructed using the flat-connection covariant derivatives ${\cal D}_{A}$. With this change of variables, we find that eq. (2.3) simplifies to
$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{{\square}}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})\exp\Big{(}\frac{{\rm i}}{\hbar}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.41)
@@
$V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{D}_{\dot{\alpha}}\Phi=0~{}.$ (2.42)
Now, repeating the analysis in [23] gives
$\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{\square}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})=1~{},$ (2.43)
and we obtain the in-out vacuum amplitude for the supersymmetric nonlinear $\sigma$-model (2.11)
$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\exp\Big{(}\frac{{\rm i}}{\hbar}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.44)
Thus we have demonstrated that the partition functions of the super FT model (2.1) and the supersymmetric nonlinear $\sigma$-model (2.11) coincide, which implies the equality of the “on-shell” effective actions in these theories[34]. This establishes the quantum equivalence of the two theories. One may also add a source term for $V$ to (2.1) and repeat the analysis to obtain formal equality of their partition functions.
The analysis in this section may be generalised by coupling the super FT model (2.1) and the supersymmetric nonlinear $\sigma$-model (2.11) to a background supergravity multiplet. Then (2.43) will turn into a topological invariant. As a result, the quantum supercurrents in the two theories $\langle J_{\alpha{\dot{\alpha}}}\rangle$ will coincide. This is an extension of the results obtained in [23, 42].
@@ 3.3 Quantisation of the free tensor multiplet
Several consistent quantisation procedures have been developed to quantise reducible Abelian gauge theories such as gauge $p$-forms [18, 19, 20, 21, 22, 23], including the formulations of [20, 23] which have been applied in the ${\cal N}=1$ supersymmetric case. These quantisation schemes are much easier to deal with than the Batalin-Vilkovisky formalism [27]. Here we will make use of the modified Faddeev-Popov procedure described in [23] to quantise the tensor multiplet model (3.6).
@@ 4 Concluding comments
$S[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\text{tr}\int{\rm d}^{4|4}z\,\mathfrak{F}(V)~{},$ (4.1)
@@
$S_{\text{T}}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\Big{\{}\chi^{\alpha}W_{\alpha}+\frac{1}{g^{2}}W^{2}\Big{\}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\Big{\{}\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{g^{2}}\bar{W}^{2}\Big{\}}~{}.$ (4.3)
@@
$\delta_{\Lambda}V=-{{\rm i}}L_{V}(\Lambda+\bar{\Lambda})+{{\rm i}}L_{V}\coth(L_{V})(\bar{\Lambda}-\Lambda)~{},\qquad\bar{D}_{\dot{\alpha}}\Lambda=0~{}.$ (4.4)
Further, it is classically topological; it describes trivial dynamics. This property proves to also hold at the quantum level, which may be seen via a slight modification of the analysis of section 2. Specifically, it is necessary to incorporate the ghost action corresponding to (4.4) and choose a gauge fermion $\Psi(\mbox{\boldmath$\Phi$})$ such that the gauge $D^{2}V=\bar{D}^{2}V=0$ is enforced. With these adjustments, a similar analysis to that conducted in section 2 leads to $Z=1$ for this theory, implying trivial dynamics.

### v2

@@ 2.2 Quantisation via Batalin-Vilkovisky formalism
$Z$ $=\int\big{[}\mathscr{D}\mbox{\boldmath$\varphi$}\big{]}\exp\Big{(}{\rm i}S[\mbox{\boldmath$\varphi$}]\Big{)}~{},\qquad\mbox{\boldmath$\varphi$}=\Big{\{}\chi^{\alpha},\bar{\chi}_{\dot{\alpha}},V,C,\tilde{C},\eta,\bar{\eta},\mbox{\boldmath$\eta$},\bar{\mbox{\boldmath$\eta$}},\eta^{\prime},\bar{\eta}^{\prime}\Big{\}}~{},$
$S[\mbox{\boldmath$\varphi$}]$ $=S_{0}[\varphi]+\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}\frac{1}{8}\tilde{C}\Big{(}\nabla^{\alpha}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}$
$\phantom{=}\quad+\frac{n}{2}\big{[}({\nabla}^{2}-4\bar{R})+(\bar{\nabla}^{2}-4R)\big{]}^{2}\Big{)}C+\frac{1}{2}(\bar{\eta}+\mbox{\boldmath$\eta$})^{2}+\frac{1}{2}(\bar{\mbox{\boldmath$\eta$}}+\eta)^{2}$
$\phantom{=}\quad+\frac{m}{2}(\bar{\eta}^{\prime}+\eta^{\prime})^{2}-\frac{m}{8}\big{(}\nabla^{\alpha}\chi_{\alpha}-\bar{\nabla}_{\dot{\alpha}}\bar{\chi}^{\dot{\alpha}}\big{)}^{2}\Big{\}}$
$\phantom{=}\quad-\frac{{\rm i}}{64}\,\text{tr}\int{\rm d}^{4|2}z_{+}\,{\cal E}\,\big{[}(\bar{\nabla}^{2}-4R)\nabla^{\alpha}\tilde{C},(\bar{\nabla}^{2}-4R)\nabla_{\alpha}\tilde{C}\big{]}\eta$
$\phantom{=}\quad+\frac{{\rm i}}{64}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{{\cal E}}\,\big{[}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\tilde{C},(\nabla^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}\tilde{C}\big{]}\bar{\eta}~{}.$ (2.30)
@@ 2.3 Quantum equivalence
$Z$ $=\int\big{[}\mathscr{D}V\big{]}\big{[}\mathscr{D}\tilde{C}\big{]}\big{[}\mathscr{D}C\big{]}\big{[}\mathscr{D}\bar{\eta}\big{]}\big{[}\mathscr{D}\eta\big{]}\big{[}\mathscr{D}\bar{\mbox{\boldmath$\eta$}}\big{]}\big{[}\mathscr{D}\mbox{\boldmath$\eta$}\big{]}\big{[}\mathscr{D}\pi\big{]}\delta\Big{(}(\nabla^{2}-4\bar{R})\pi\Big{)}\delta\Big{(}(\bar{\nabla}^{2}-4R)\pi\Big{)}$
×$\qquad\qquad\times\delta\Big{(}W_{\alpha}+\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4{R})\nabla_{\alpha}\pi\Big{)}\delta\Big{(}\bar{W}_{\dot{\alpha}}-\frac{{\rm i}}{4}(\nabla^{2}-4\bar{R})\bar{\nabla}_{\dot{\alpha}}\pi\Big{)}\text{Det}(\mathcal{H})$
×$\qquad\qquad\times\exp\Big{(}{\rm i}\,\text{tr}\int{\rm d}^{4|4}z\,E\,\Big{\{}V^{2}-2\tilde{C}\Box_{V}C+\frac{1}{2}({\mbox{\boldmath$\eta$}}+\bar{\eta})^{2}+\frac{1}{2}({\eta}+\bar{\mbox{\boldmath$\eta$}})^{2}\Big{\}}\Big{)}~{},$ (2.31)
@@
$\mathcal{H}$ $=\begin{pmatrix}R&-\frac{1}{4}(\bar{\nabla}^{2}-4R)\\ -\frac{1}{4}({\nabla}^{2}-4\bar{R})&\bar{R}\end{pmatrix}~{},$ (2.32a)
$\Box_{V}$ $=-\frac{1}{16}\Big{(}\nabla^{\alpha}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}+\bar{\nabla}_{\dot{\alpha}}({\nabla}^{2}-4\bar{R})\bar{\nabla}^{\dot{\alpha}}-\big{[}({\nabla}^{2}-4\bar{R})+(\bar{\nabla}^{2}-4R)\big{]}^{2}\Big{)}$
$=\nabla^{a}\nabla_{a}-\frac{1}{4}G^{{\alpha{\dot{\alpha}}}}[\nabla_{\alpha},\nabla_{\dot{\alpha}}]-\frac{1}{4}\bar{\nabla}_{\dot{\alpha}}\bar{R}\bar{\nabla}^{\dot{\alpha}}-\frac{1}{4}\nabla^{\alpha}R\nabla_{\alpha}-\frac{1}{4}\bar{\nabla}^{2}\bar{R}-\frac{1}{4}\nabla^{2}R$
$\phantom{=}+2R\bar{R}-\frac{1}{4}R(\bar{\nabla}^{2}-4R)-\frac{1}{4}\bar{R}({\nabla}^{2}-4\bar{R})-\frac{1}{4}W^{\alpha}\nabla_{\alpha}+\frac{1}{4}\bar{W}_{\dot{\alpha}}\bar{\nabla}^{\dot{\alpha}}~{}.$ (2.32b)
The operator $\mathcal{H}$ acts on the space of covariantly chiral-antichiral pairs in the adjoint representation of $G$, while $\Box_{V}$ is defined to act on the space of unconstrained scalar superfields in the adjoint representation of $G$. The operators $\mathcal{H}$ and $\Box_{V}$ are gauge-covariant extensions of the operators introduced in [51].
@@
$W_{\alpha}={\rm e}^{\bar{\Omega}_{Q}}\left(-\frac{1}{4}(\bar{\cal D}^{2}-4{R})\Big{(}{\rm e}^{-V_{Q}}{\cal D}_{\alpha}{\rm e}^{V_{Q}}\Big{)}+w_{\alpha}\right){\rm e}^{-\bar{\Omega}_{Q}}~{},$ (2.35)
@@
$W_{\alpha}\approx-\frac{1}{4}(\bar{\cal D}^{2}-4{R}){\cal D}_{\alpha}V_{Q}+w_{\alpha}+\big{[}\bar{\Omega}_{Q},w_{\alpha}\big{]}~{}.$ (2.36)
@@
$w_{\alpha}=0\quad\implies\quad{\rm e}^{\bar{\omega}}={\rm e}^{\Phi}~{},\qquad{\bar{\mathfrak{D}}}_{\dot{\alpha}}\Phi=0~{}.$ (2.37)
In evaluating $\delta(W_{\alpha}+\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4{R})\nabla_{\alpha}\pi)$, it suffices to keep only the terms of first order in the quantum prepotentials and $\pi$,
$\delta\left(W_{\alpha}+\frac{{\rm i}}{4}(\bar{\nabla}^{2}-4R)\nabla_{\alpha}\pi\right)=\delta\left(-\frac{1}{4}(\bar{{\cal D}}^{2}-4{R}){\cal D}_{\alpha}V_{Q}+\frac{{\rm i}}{4}(\bar{{\cal D}}^{2}-4{R}){\cal D}_{\alpha}\pi\right)~{}.$ (2.38)
@@
$V_{Q}$ $=\rho+\bar{\rho}+\frac{1}{2}\big{(}{\cal D}^{\alpha}\lambda_{\alpha}+\bar{{\cal D}}_{\dot{\alpha}}\bar{\lambda}^{\dot{\alpha}}\big{)}~{},\qquad\bar{{\cal D}}_{\dot{\alpha}}\rho=0~{},\quad{\bar{\cal D}}_{\dot{\alpha}}\lambda_{\alpha}=0~{},$ (2.39a)
@@
$1=\int\big{[}\mathscr{D}V_{Q}\big{]}\big{[}\mathscr{D}\pi\big{]}\exp\Big{(}\frac{{\rm i}}{2}\text{tr}\int\,{\rm d}^{4|4}z\,E\,\Big{\{}{V_{Q}}^{2}-\pi^{2}\Big{\}}\Big{)}$ (2.40)
@@
~$\widetilde{\mathscr{H}}_{c}=\frac{1}{16}(\bar{{\cal D}}^{2}-4{R})({\cal D}^{2}-6\bar{R})~{},\qquad\widetilde{\mathscr{H}}_{a}=\frac{1}{16}({\cal D}^{2}-4\bar{{R}})(\bar{{\cal D}}^{2}-6{{R}})~{}.$ (2.42)
Here and below, a tilde over operators, such as $\widetilde{{\cal H}}$, indicates that these operators are constructed using the flat-connection covariant derivatives ${\cal D}_{A}$. With this change of variables, we find that eq. (2.3) simplifies to
$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{{\square}}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})\exp\Big{(}{\rm i}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.43)
@@
$V=\log({\rm e}^{\bar{\Phi}}{\rm e}^{\Phi}\big{)}~{},\qquad\bar{\mathfrak{D}}_{\dot{\alpha}}\Phi=0~{}.$ (2.44)
Now, the analysis in [23] (see also [44] for a review) implies that
$\text{Det}_{+}^{1/2}(\widetilde{\mathscr{H}}_{c})\text{Det}_{-}^{1/2}(\widetilde{\mathscr{H}}_{a})\text{Det}(\widetilde{\square}_{V})\text{Det}^{-2}(\widetilde{\mathcal{H}})~{},$ (2.45)
is a topological invariant, that is it does not depend on the supergravity prepotentials $H^{m}$, $\varphi$ and $\bar{\varphi}$, where $H^{m}$ is the gravitational superfield [53, 54] and $\varphi$ the so-called chiral compensator [54] (see [44, 47] for a review). Therefore, the in-out vacuum amplitude (2.43) reduces to that for the supersymmetric nonlinear $\sigma$-model (2.13)
$Z$ $=\int\big{[}\mathscr{D}\bar{\Phi}\big{]}\big{[}\mathscr{D}{\Phi}\big{]}\exp\Big{(}{\rm i}S[\Phi,\bar{\Phi}]\Big{)}~{}.$ (2.46)
Thus we have demonstrated that the partition functions of the super FT model (2.1) and the supersymmetric nonlinear $\sigma$-model (2.13) coincide, which implies the equality of the “on-shell” effective actions in these theories[34]. This establishes the quantum equivalence of the two theories in the sense that the quantum supercurrents of the two theories coincide. One may also add a source term for $V$ to (2.1) and repeat the analysis to obtain formal equality of their partition functions. This is an extension of the results obtained in [23, 42].
@@ 3.3 Quantisation of the free tensor multiplet
Several consistent quantisation procedures have been developed to quantise reducible Abelian gauge theories such as gauge $p$-forms [18, 19, 20, 21, 22, 23], including the formulations of [20, 23] which have been applied in the ${\cal N}=1$ supersymmetric case. These quantisation schemes are much easier to deal with than the Batalin-Vilkovisky formalism [27]. Here we will make use of the modified Faddeev-Popov procedure described in [23] to quantise the tensor multiplet model (3.6). [^10] Quantisation of this model within the BV formalism is described in appendix B.
@@ 4 Concluding comments
$S[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\mathcal{E}\,\chi^{\alpha}W_{\alpha}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\mathcal{E}}\,\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\text{tr}\int{\rm d}^{4|4}z\,E\,\mathfrak{F}(V)~{},$ (4.1)
@@
$S_{\text{T}}[\chi,\bar{\chi},V]=\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{+}\,\mathcal{E}\,\Big{\{}\chi^{\alpha}W_{\alpha}+\frac{1}{g^{2}}W^{2}\Big{\}}+\frac{1}{2}\text{tr}\int{\rm d}^{4|2}z_{-}\,\bar{\mathcal{E}}\,\Big{\{}\bar{\chi}_{\dot{\alpha}}\bar{W}^{\dot{\alpha}}+\frac{1}{g^{2}}\bar{W}^{2}\Big{\}}~{}.$ (4.3)
@@
$\delta_{\Lambda}V=-{{\rm i}}L_{V}(\Lambda+\bar{\Lambda})+{{\rm i}}L_{V}\coth(L_{V})(\bar{\Lambda}-\Lambda)~{},\qquad\bar{\mathfrak{D}}_{\dot{\alpha}}\Lambda=0~{}.$ (4.4)
Further, it is classically topological; it describes trivial dynamics. This property proves to also hold at the quantum level, which may be seen via a slight modification of the analysis of section 2. Specifically, it is necessary to incorporate the ghost action corresponding to (4.4) and choose a gauge fermion $\Psi(\mbox{\boldmath$\Phi$})$ such that the gauge $(\mathfrak{D}^{2}-4\bar{R})V=(\bar{\mathfrak{D}}^{2}-4R)V=0$ is enforced. With these adjustments, a similar analysis to that conducted in section 2 leads to $Z=1$ for this theory, implying trivial dynamics.
@@
An ${\cal N}=2$ locally supersymmetric analogue of the topological model (4.3) is
$S=\frac{1}{2}\int{\rm d}^{4}x{\rm d}^{4}\theta\,{\cal E}\,W\big{(}\Psi+\frac{1}{g^{2}}W\big{)}+{\rm c.c.}~{},$ (4.9)
where the prepotential $\Psi$ for the tensor multiplet is now covariantly chiral, and chiral field strength $W$ for the vector multiplet obeys the constraints
${\bar{\cal D}}^{\dot{\alpha}}_{i}W=0~{},\qquad\Big{(}{\cal D}^{\alpha(i}{\cal D}_{\alpha}^{j)}+4S^{ij}\Big{)}W=\Big{(}{\bar{\cal D}}_{\dot{\alpha}}{}^{(i}{\bar{\cal D}}^{j){\dot{\alpha}}}+4\bar{S}^{ij}\Big{)}\bar{W}~{}.$ (4.10)
This model respects the gauge symmetries of both tensor and vector multiplets. It would be interesting to compute the partition function for the model since its logarithmically divergent part should contain the ${\cal N}=2$ Gauss-Bonnet term [74].